OpenStudy (anonymous):

find F'(x) when F(x)= integral from 4 to x^2 (2 sqrt(1+t^3) dt)

OpenStudy (freckles):

\[F(x)=\int\limits_4^{x^2} g(t) dt=G(t)|_4^{x^2}=G(x^2)-G(4) \\ \text{so we have } F(x)=G(x^2)-G(4) \\ \text{ differentiate both sides }\]

OpenStudy (freckles):

\[F'(x)=(x^2)'g(x^2)-0 \text{ by chain rule and constant rule }\]

OpenStudy (freckles):

@aawowaa are you there?

OpenStudy (irishboy123):

+

OpenStudy (anonymous):

yes, I'm going over it.

OpenStudy (freckles):

you understand that g(t) is 2sqrt(1+t^3)

OpenStudy (freckles):

And that I was using G(t) to represent the antiderivative of g(t) that is G'=g

OpenStudy (anonymous):

yes, I understand

OpenStudy (anonymous):

so now I plug in 2sqrt(1+t^3) back?

OpenStudy (freckles):

g(t)=2sqrt(1+t^3) so g(x^2)=?"

OpenStudy (anonymous):

=2sqrt(1+x^6)

OpenStudy (freckles):

ok so that is what you replace g(x^2) with

OpenStudy (freckles):

don't forget to find (x^2)' which is next to g(x^2)

OpenStudy (anonymous):

thank you I got it

OpenStudy (freckles):

np